OpenStudy (anonymous):

If i is raised to an even power, then it can not simplify to be -1 1 i

hartnn (hartnn):

do you know the definition of 'i' ?

hartnn (hartnn):

find i^2 i^4 i^6 i^8 ... see the pattern, you'll get your answer :)

OpenStudy (anonymous):

yes its the imaginary quantity equal to the square root of minus one.

OpenStudy (anonymous):

I know the definition, not how to do it.

hartnn (hartnn):

so \(i = \sqrt{-1}\) square it, what do u get ?

OpenStudy (anonymous):

I don't understand the entire content of "imaginary numbers" it doesn't click with me. How can I answer it if i don't understand it?

hartnn (hartnn):

you understand squaring ? and \((\sqrt{a})^2 =a\) thats basic exponent rule. if you're not comfortable with "imaginary" term, assume 'i' as a variable, like x

OpenStudy (anonymous):

I have tried everything to learn imaginary numbers. teachers, tutors, anything to help my pass. I don't have a calculator so squaring by hand is a pain. let alone imaginary numbers. My brain just doesn't want to understand it. it makes no sense what so ever.

hartnn (hartnn):

if \(x = \sqrt a\) so, \(x^2 =a\) do u understand that ? nothing to do with imaginary numbers...

OpenStudy (anonymous):

kind of.

hartnn (hartnn):

ok, then try to answer this : if \(i = \sqrt{-1} \\ then, \: i^2 = .. ?\)

OpenStudy (anonymous):

1?

hartnn (hartnn):

why 1? and why not -1 ? where did the negative go ? did you eat it ? :P

OpenStudy (anonymous):

could have swore i put -1. I'm on a crap cromebook thinkpad thing, spacebar halfway works.

OpenStudy (anonymous):

yes, I got hungry so i ate the negative.

hartnn (hartnn):

cool, so, \(i^2 =-1\) is correct :) now square both sides again, what do u get ?

OpenStudy (anonymous):

1

OpenStudy (anonymous):

I think

hartnn (hartnn):

correct again! :)

hartnn (hartnn):

so, i^2 =-1 i^4 = 1 so which value you cannot get ? out of 1,-1, i ?

OpenStudy (anonymous):

i

hartnn (hartnn):

thats your answer :)

hartnn (hartnn):

please refer to this also, maybe you'll start understanding things :) http://www.mathsisfun.com/numbers/imaginary-numbers.html

hartnn (hartnn):

feel free to ask any doubts :)

OpenStudy (anonymous):

thank you!

hartnn (hartnn):

welcome ^_^